OpenStudy (anonymous):

Suppose that \( b^2 - 4ac < 0\). Show that there are no numbers x satisfying\[x^2 + bx + c = 0\]In fact, \( x^2 + bc + c > 0\) for all x.

OpenStudy (jamesj):

where's your a gone? And in fact, it could be strictly positive or strictly negative.

OpenStudy (anonymous):

Imaginary numbers, becuase u have a square root of a negative.

OpenStudy (anonymous):

I don't know where the a went :( that's the question out of the book

OpenStudy (anonymous):

it says hint: complete the square

OpenStudy (anonymous):

is this of diff eqs??? becuase i just took a diff eqs test and it had this on it, so imaginary number are the product of b2−4ac<0

OpenStudy (anonymous):

It's basic calculus

OpenStudy (anonymous):

from Spivak

OpenStudy (jamesj):

Well if we assume that a = 1, then \[ x^2 + bx + c = (x + b/2a)^2 + 1/4a^2 ( 4ac - b^2 ) > 0 \ \ \ \ \forall \ x \]

OpenStudy (anonymous):

Now Spivak asks to prove that \( x^2 + xy + y^2 > 0\) provided that x and y are not both 0.

OpenStudy (anonymous):

I'm curious about that one. We know that x^2 and y^2 will always be positive regardless of the sign of x or y. I wonder how would can verify that x^2 + y^2 > xy, for x or y being negative.

OpenStudy (jamesj):

Well, it's fairly obvious what he wants you to do. EITHER treat this as a quadratic in x or y and use the relation you just proved. OR repeat this method of completing the square

OpenStudy (anonymous):

Ah ha. \[\left ( x + {y \over 2} \right)^2 + {2y^2 \over 4}\] All quantities are squared. http://www.wolframalpha.com/input/?i=x%5E2+%2B+xy+%2B+y%5E2

OpenStudy (anonymous):

Rather 3/4 in the second term.